A man walks a distance of 3 units from the origin towards the north-east( N 45 degrees E) direction. From there he walks a distance of 4 units towards the north-west ( N 45 degrees W) direction to reach a point P.Then the position of P in the Argand plane is:

 

Ans. given is (3+4i)e^(i.pi/4)

 

Grade:12th pass

1 Answers

Vikas TU
7 years ago
If z is the complex number then,
(z – 3e^i*pi/4)/(0 – 3*e^ipi/4) = (4/3)*e^-i*pi/2 = -4i/3
solving we get,
z = (3 + 4i)*e^ipi/4